At the present time Albert is 20 years old and Steven is 16 years old. How many years ago was Albert 1 1/2 times as old as Steven?
Answer by macston(5194)

.
A=Albert's age; S=Steven's age; x=years ago
.
A-x=1.5(S-x)
20-x=1.5(16-x)
20-x=24-1.5x
0.5x=4
x=8
ANSWER Albert was 1.5 times as old as Steven 8 years ago.